40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Nov 13 22:32:54 2015, Koji, Summary, PSL, PMC LO degraded, usual ERA-5 replacement, LO recovered PMC_LO.pngIMG_2093.JPGIMG_2091.JPGIMG_2095.JPGIMG_2096.JPG
    Reply  Sun Nov 15 22:43:48 2015, Koji, Summary, PSL, PMC LO degraded, usual ERA-5 replacement, LO recovered C1-MULTI_E6875C_TIMESERIES-1131408017-86400.pngC1-MULTI_E6875C_TIMESERIES-1131580817-86400.png
       Reply  Mon Nov 16 19:05:59 2015, Koji, Summary, PSL, PMC servo circuit review, follow up measurements 
          Reply  Mon Nov 16 21:32:49 2015, Koji, Summary, PSL, PMC servo circuit review, follow up measurements 
             Reply  Tue Nov 17 16:21:10 2015, Koji, Summary, PSL, PMC servo circuit review, follow up measurements 
                Reply  Wed Nov 18 16:51:58 2015, Koji, Summary, PSL, PMC servo calibration PMC_err_cal.pdf
                   Reply  Wed Nov 18 16:53:17 2015, Koji, Summary, PSL, PMC PZT capacitance 
                      Reply  Wed Nov 18 17:09:22 2015, Koji, Summary, PSL, PMC Servo analysis PMC_OLTF.pdfPMC_DSP.pdf
                         Reply  Fri Nov 20 15:44:12 2015, Koji, Summary, PSL, Added 17.5kHz LPF to the PMC servo PMC_OLTF.pdfPMC_noise_comparison.pdf
Message ID: 11769     Entry time: Mon Nov 16 21:32:49 2015     In reply to: 11768     Reply to this: 11775
Author: Koji 
Type: Summary 
Category: PSL 
Subject: PMC servo circuit review, follow up measurements 

The result of the precise inspection for the PMC servo board for the 40m was done.

The record, including the photo of the board, can be found at https://dcc.ligo.org/D1400221-v2

- I found some ceramic 1uF caps are used in the signal path. They have been replaced with film caps by WIMA.

- In later measurements with the openloop TF measurement, it was found that the notch frequency (14.6kHz) was off from the a sharp PZT resonance at 12.2kHz.
I replaced the combined caps of 1220pF to 1742pF. This resulted nice agreement of the notch freq with the PZT resonant freq.


Past related elogs:

SRA-3MH mixer installed in 2009: http://nodus.ligo.caltech.edu:8080/40m/1502

R20 increased for more LO Mon gain: http://nodus.ligo.caltech.edu:8080/40m/10172

ELOG V3.1.3-